# TiDES: The 4MOST Time Domain Extragalactic Survey

### v1

@@ 3 Simulating LSST+TiDES
We next describe our method for simulating a LSST-like SN survey, and combining with a simulated 4MOST observing strategy. Our simulations focus on LSST for the transient discovery as we anticipate this will provide the majority of the TiDES targets. However, TiDES will be agnostic to the source of the photometric transient feed and will include transients from any source providing a public alert stream (e.g., Bellm et al., 2019; Jones et al., 2021; Kim et al., 2022; Groot et al., 2022). Thus, this section is intended to produce broad projections for the number and type of transients and hosts that 4MOST may observe, in particular noting that neither LSST nor 4MOST’s strategy are yet finalised. An end-to-end flowchart of our entire simulation framework is shown in Fig. 1.
@@ 3.1 Simulating transients
The LSST Science Requirements Document (Ivezić & the LSST Science Collaboration, 2013) details the survey specification but with enough room for optimisation of the strategy to balance the different science goals (Bianco et al., 2022). LSST will have a ‘Wide, Fast, Deep’ (WFD) component coupled with deeper, more frequent observations in ‘Deep Drilling Fields’ (DDFs). Significant work has been performed on the effects of different survey choices for transients (e.g., Lochner et al., 2022; Alves et al., 2023; Gris et al., 2023), and we do not attempt to determine an optimal LSST strategy for TiDES in this work. Instead, we use established LSST simulations to create our mock transient datasets, adopting the baseline v3.4111https://survey-strategy.lsst.io/index.html simulation, the latest baseline release at the time of writing. The comparisons with other LSST strategies are presented in the TiDES Cadence Note222 https://docushare.lsst.org/docushare/dsweb/Get/Document-37640/Frohmeier_TiDES.pdf. Note that due to the high cadence of repeat visits in the DDFs, all of our TiDES-RM targets are located there.
The output from the LSST Operations Simulator (OpSim)333https://www.lsst.org/scientists/simulations/opsim is the input to our simulations. Based on a scheduling pattern, OpSim creates a catalogue-level output of the LSST strategy using detailed observing quality metrics based on weather models, maintenance requirements, and telescope hardware constraints. Using these observing characteristics, we then simulate transients and their host galaxies using the SuperNova ANAlysis software snana (Kessler et al., 2009) simulation code. snana combines survey-specific metadata (e.g., from OpSim) with the known properties of a SN population to generate a set of light curves as observed by any survey strategy. The OpSim strategies are translated into the appropriate snana format using the OpSimSummary package (Biswas et al., 2020). Note that all of our simulations are ‘catalogue level’, i.e., we do not place transients and hosts into simulated images.
snana generates light curves by integrating a time-evolving spectral energy distribution (‘template’) of a given SN through the survey’s filter-specific response function. Table LABEL:tab:snClasses shows each spectral template we use in our simulations. Our method of generating light curves from spectral templates using snana follows the Photometric LSST Astronomical Time Series Classification Challenge (PLAsTiCC) (Kessler et al., 2019a), but we update the core collapse SN templates to those of Vincenzi et al. (2019, hereafter V19).
@@
SN Ia | Normal type Ia SNe following the SALT2 model | Guy et al. (2007); Hounsell et al. (2018) |
SN Iax | Type Iax peculiar SNe | Kessler et al. (2019a) |
SN Ia-91bg | Peculiar SNe SN1991bg-like | Kessler et al. (2019a) |
Type II core collapse (SN II, IIn) | Hydrogen-rich core-collapse SNe | Vincenzi et al. (2019) |
Stripped Envelope Supernovae (SN IIb, Ib, Ic, Ic-BL) | Hydrogen-poor stripped envelope SNe | Vincenzi et al. (2019) |
CaRT/CaST | Calcium-rich(-strong) faint and fast SNe | Kessler et al. (2019a) |
SLSN-I | Hydrogen-poor super luminous SNe | Nicholl et al. (2017); Kessler et al. (2019a) |
TDE | Tidal disruption events | Kessler et al. (2019a); Mockler et al. (2019) |
@@ 3.1.1 Simulation of SNe Ia
We simulate SNe Ia using the SALT2 SED model (Guy et al., 2007, 2010) trained following Betoule et al. (2014). The SALT2 SED model is described by the SN redshift , the epoch of SN peak brightness , a (dimensionless) stretch-like parameter , and a colour parameter . The SN Ia apparent magnitude in the rest-frame -band is then simulated as
where the parameters and describe the stretch–luminosity and colour–luminosity relations and are set in our simulations to , and . The absolute magnitude for a , SN Ia () is set to -19.365 and is the luminosity distance for redshift with our default cosmological parameters. The term is a randomised brightness offset that encapsulates the intrinsic scatter observed in SN Ia standardised brightness (see Guy et al., 2010, for details on the modelling of intrinsic scatter). The absolute number of SN Ia events simulated and their redshift distribution is calculated using the power-law evolution of the volumetric rates presented by Frohmaier et al. (2019) (their Equation 6). The intrinsic distributions of SN stretch and colour are taken from Scolnic & Kessler (2016).
@@ 3.1.2 Simulation of other types of transients
All simulated transient types are listed in Table LABEL:tab:snClasses. For core-collapse SNe, we use V19 templates and include both hydrogen-rich core-collapse SN events (type II and IIn SNe) and stripped envelope SNe (Type Ib/IIb/Ic/Ic-BL SNe). For volumetric rates of core-collapse events, we normalise at low redshift to the measurement of Frohmaier et al. (2021) and extrapolate to higher redshifts assuming the rate follows the star formation history presented in Madau & Dickinson (2014). To ensure the correct mix of different core collapse sub-types we use the relative rates presented by Shivvers et al. (2017). The intrinsic brightness of simulated events are matched to the luminosity functions presented in Li et al. (2011) and revised by V19. Recently, updates have been made to models using different rate assumptions for Type IIn SNe (Ransome & Villar, 2024) and templates for stripped-envelope (Khakpash et al., 2024), however, these updates do not affect the overall picture for TiDES compared to the models we have used.
We include two types of peculiar SNe Ia: SN1991bg-like events (Filippenko et al., 1992) and SNe Iax (Foley et al., 2013). SN1991T-like events are modelled within the normal-SN Ia events in Section 3.1.1. The SED templates and volumetric rates used to simulate these classes of peculiar SNe Ia are those presented by Kessler et al. (2019a), applying the revisions in Vincenzi et al. (2021, hereafter V21).
Tidal Disruption Events (TDEs), Calcium-Strong Transients (CaST) and hydrogen-poor Superluminous SNe (SLSNe-I) are implemented following the same assumptions and modelling used in the PLASTICC challenge (Kessler et al., 2019a).
@@ 3.1.3 Simulation of AGN for TiDES-RM
The AGN that we will observe as part of the reverberation mapping survey are already known and are located only in the LSST DDFs. The selection is based on two catalogues, the Million Quasar Catalogue Milliquas (Flesch, 2023) and the southern photometric quasar catalogue based on Dark Energy Survey DR2 as presented by Yang & Shen (2023). First, both catalogues were cut down to the hexagonal footprints of the 4MOST field-of-view centered on the defined 4MOST deep field positions within the Elais South, COSMOS, ECDF-S, and XMM-LSS deep fields (Table LABEL:tab:ddf_table). For Milliquas, only sources with a Q or A type flag have been included, which selects against obscured AGN that are unfavourable to reverberation mapping. In a second step, sources from both catalogues were cross-matched. Those present in both catalogues are deemed the “gold sample” while objects only present in Milliquas are part of the “silver sample.” The total number of identified AGN varies with deep field, given how intensely the respective field has been monitored and in which wavebands. Given survey operational constraints, the maximum number of TiDES-RM sources that can be observed per deep field is of the order of 250. To preserve some comparability across the fields, we impose brightness cuts of for the COSMOS and XMM-LSS fields and for the Elais South and ECDF-S fields.
DDF | R.A. | Dec. |
ELAIS-S1 | 9.500 | -43.950 |
XMM-LSS | 35.875 | -5.025 |
ECDF-S | 53.125 | -28.100 |
COSMOS | 150.125 | 2.200 |
@@
Finally, we want to exclude sources where the survey length will not be able to recover a lag. This requires estimating observed lag length of the cut down catalogues. We use the lag-luminosity relation for the H line presented in Bentz et al. (2009) and scale to Mg ii lags with a factor of 0.9. C iv lags are estimated from the lag-luminosity relation in Penton et al. (2022). Approximate luminosities are calculated from band magnitudes. The resulting estimated rest-frame lags have been converted to observed lags by accounting for a cosmological time dilation factor of . We assume that successful lag recovery requires lag lengths to be at least 3 times shorter than the survey length. As such, we cut the target catalogue for AGN where the shortest lag observed in the 4MOST wavelength range for the given redshift is 600 days. With the previously imposed brightness limit, this primarily affects the brightest sources in the redshift range and particularly those at the higher redshift end, where Mg ii is the only one of the 3 main target lines present. For all four DDFs, this affects 1% of the pre-cut samples.
Field | total | H | Mg ii | C iv |
ELAIS-S1 | 191 | 23 | 128 | 80 |
XMM-LSS | 252 | 47 | 141 | 119 |
ECDF-S | 230 | 48 | 138 | 92 |
COSMOS | 256 | 47 | 150 | 94 |
@@
Notes — Redshift ranges covered by indicidual lines are: H: ; MgII : ; C iv: . Spectra of some sources will include two broad lines that can be used to cross-calibrate lags: ELAIS-S1: H/Mg ii: 10, Mg ii/C iv: 30; XMM-LSS: H/Mg ii: 17, Mg ii/C iv: 38; ECDF-S: H/Mg ii: 20, Mg ii/C iv: 28; COSMOS: H/Mg ii: 14, Mg ii/C iv: 21.
@@ 3.2 Simulating transient host galaxies
SNe Ia, peculiar SNe Ia and core-collapse SNe are associated to host galaxies following the approach and assumptions presented in V21. Host galaxies are selected from a ‘HOSTLIB’ library of 380 000 galaxies from the DES Science Verification galaxy catalogues (Drlica-Wagner et al., 2018). Each galaxy in the catalogue has optical photometry from DES () and – when available – near infrared photometry from VISTA. To derive properties of the galaxies, template SEDs were generated using the PÉGASE.2 spectral synthesis code and fit to the observed galaxy fluxes through a minimisation; this routine is extensively described in Sullivan et al. (2010) and, in the context of DES galaxy catalogues, Smith et al. (2020). From the best fitting templates, we adopt the associated parameters such as: redshift, stellar mass, and star formation rates throughout the rest of this work. The catalogue is complete to mag in -band, which is well-matched to the transient hosts that LSST should detect (e.g., Wiseman et al., 2020).
Host galaxies are associated to simulated SNe Ia following the observed rates as a function of the host-galaxy properties as measured by Sullivan et al. (2006). Within each galaxy, we follow the standard snana prescription of placing transients in the host following a Sérsic profile (Sérsic, 1963). We also ensure that the well-known correlation between the SN Ia parameter and host galaxy stellar mass is reproduced in the simulations (see section 4.5.1 in V21 for details), but we do not include the more subtle relation between SN Ia Hubble residuals and host galaxy properties, such as stellar mass, colour, and specific star formation rate.
For peculiar SNe Ia, we scale the SN Ia rate and assign host galaxies so that SN1991bg-like events are only associated to old stellar environments (e.g. Li et al., 2011; Panther et al., 2019) and SN Iax events are placed in star forming galaxies (e.g. Lyman et al., 2013, 2018; Takaro et al., 2020). For core-collapse SNe, we use the rates and host galaxy association measurements presented by Graur et al. (2017a, b). For the remaining transients their relation to host galaxies is not as extensively studied due to their rarer intrinsic and observational rates; we assign them to hosts randomly and position them within the host following a Sérsic profile. This random host association for the sub-dominant populations has no significant effect on the TiDES host galaxy statistics.
@@ 3.3 Simulating AGN
AGN simulations have been carried out based on properties of the selected sample. Due to the nature of optical selection in the target catalogues, the overwhelming majority of the sources will be broad line AGN, as desired for a reverberation mapping experiment. We identify each selected source with a spectroscopic template of a type 1 AGN as provided by the 4MOST AGN survey (Merloni et al., 2019). The associated template spectrum for each object has been redshifted and flux scaled according to the catalogue -band magnitude and fed into the 4MOST operations simulator. Within a 4MOST simulation, the AGN templates are used to determine the required exposure times in the DDFs to meet the spectral success criteria (Section 4.2).
@@ 3.4 Simulating 4MOST
Having generated a simulation of transient light curves as observed by LSST, we next determine which are available to be observed by 4MOST. The 4MOST observing strategy fulfils a complex balance of science requirements from multiple different surveys focused on both galactic and extra-galactic fields. To accommodate this, 4MOST will tile the entire southern-sky over the course of its five-year lifetime (de Jong et al., 2019), with the density of targets and exposure time requirements ultimately dictating the frequency of any returning visits. The overlap between the LSST wide-fast-deep footprint and the 4MOST union of observing tiles is shown in Fig. 2.
@@
4MOST has experimented with several mock strategies to establish its observing plan. The success of any strategy is evaluated using a figure-of-merit for each 4MOST survey, combined into a global figure of merit. We evaluate the success of TiDES using the total number of objects observed and the quality of their resulting spectra. Our simulations in this paper use a typical 4MOST mock strategy that has been optimised to a solution that tiles the sky with observations (Tempel et al., 2020a) using probabilistic fibre assignments (Tempel et al., 2020b). From the TiDES perspective, we can use these simulations to track which of our simulated transients were observed and the exposure time on that epoch. For the host galaxies, we are able to assess the total integrated time the target received over multiple epochs.

### v2

@@ 3 Simulating LSST+TiDES
We next describe our method for simulating a LSST-like SN survey, and combining with a simulated 4MOST observing strategy. Our simulations focus on LSST for the transient discovery as we anticipate this will provide the majority of the TiDES targets. However, TiDES will be agnostic to the source of the photometric transient feed and will include transients from any source providing a public alert stream e.g. the Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019), the Young Supernova Experiment (YSE; D. O. Jones et al., 2021), BlackGEM (P. J. Groot et al., 2022), The Gravitational-wave Optical Transient Observer (GOTO; D. Steeghs et al., 2022), and The La Silla Schmidt Southern Survey (LS4; A. A. Miller et al., 2025). Thus, this section is intended to produce broad projections for the number and type of transients and hosts that 4MOST may observe, in particular noting that neither LSST nor 4MOST’s strategy are yet finalised. An end-to-end flowchart of our entire simulation framework is shown in Fig. 1.
@@ 3.1 Simulating transients
| SN Ia | Normal type Ia SNe following SALT2 | R. Hounsell et al. (2018) |
| SN Iax | Type Iax peculiar SNe | R. Kessler et al. (2019a) |
| SN Ia-91bg | Peculiar SNe SN1991bg-like | R. Kessler et al. (2019a) |
| SN II, IIn | Hydrogen-rich core-collapse SNe | M. Vincenzi et al. (2019) |
| SN IIb, Ib, Ic, Ic-BL | Hydrogen-poor stripped envelope SNe | M. Vincenzi et al. (2019) |
| CaRT/CaST | Calcium-rich(-strong) faint and fast SNe | R. Kessler et al. (2019a) |
| SLSN-I | Hydrogen-poor super luminous SNe | M. Nicholl et al. (2017); R. Kessler et al. (2019a) |
| TDE | Tidal disruption events | R. Kessler et al. (2019a); B. Mockler et al. (2019) |
The LSST Science Requirements Document (Ž.. Ivezić & the LSST Science Collaboration, 2013) details the survey specification but with enough room for optimisation of the strategy to balance the different science goals (F. B. Bianco et al., 2022). LSST will have a ‘Wide, Fast, Deep’ (WFD) component coupled with deeper, more frequent observations in ‘Deep Drilling Fields’ (DDFs). Significant work has been performed on the effects of different survey choices for transients (e.g., M. Lochner et al., 2022; C. S. Alves et al., 2023; P. Gris et al., 2023), and we do not attempt to determine an optimal LSST strategy for TiDES in this work. Instead, we use established LSST simulations to create our mock transient datasets, adopting the baseline v3.4111https://survey-strategy.lsst.io/index.html simulation, the latest baseline release at the time of writing. The comparisons with other LSST strategies are presented in the TiDES Cadence Note222 https://docushare.lsst.org/docushare/dsweb/Get/Document-37640/Frohmeier_TiDES.pdf. Note that due to the high cadence of repeat visits in the DDFs, all of our TiDES-RM targets are located there.
The output from the LSST Operations Simulator (OpSim)333https://www.lsst.org/scientists/simulations/opsim is the input to our simulations. Based on a scheduling pattern, OpSim creates a catalogue-level output of the LSST strategy using detailed observing quality metrics based on weather models, maintenance requirements, and telescope hardware constraints. Using these observing characteristics, we then simulate transients and their host galaxies using the SuperNova ANAlysis software snana (R. Kessler et al., 2009) simulation code. snana combines survey-specific metadata (e.g., from OpSim) with the known properties of a SN population to generate a set of light curves as observed by any survey strategy. The OpSim strategies are translated into the appropriate snana format using the OpSimSummary package (R. Biswas et al., 2020). Note that all of our simulations are ‘catalogue level’, i.e., we do not place transients and hosts into simulated images. We assume a saturation limit of 15 mag for the LSST CCD and exclude all points on a simulated light curve exceeding this brightness.
snana generates light curves by integrating a time-evolving spectral energy distribution (‘template’) of a given SN through the survey’s filter-specific response function. Table LABEL:tab:snClasses shows each spectral template we use in our simulations. Our method of generating light curves from spectral templates using snana follows the Photometric LSST Astronomical Time Series Classification Challenge (PLAsTiCC) (R. Kessler et al., 2019a), but we update the core collapse SN templates to those of M. Vincenzi et al. (2019, hereafter V19).
@@ 3.1.1 Simulation of SNe Ia
We simulate SNe Ia using the SALT2 SED model (J. Guy et al., 2007, 2010) trained following M. Betoule et al. (2014). The SALT2 SED model is described by the SN redshift , the epoch of SN peak brightness , a (dimensionless) stretch-like parameter , and a colour parameter . The SN Ia apparent magnitude in the rest-frame -band is then simulated as
where the parameters and describe the stretch–luminosity and colour–luminosity relations and are set in our simulations to , and . The absolute magnitude for a , SN Ia () is set to -19.365 and is the luminosity distance for redshift with our default cosmological parameters. The term is a randomised brightness offset that encapsulates the intrinsic scatter observed in SN Ia standardised brightness (see J. Guy et al., 2010, for details on the modelling of intrinsic scatter). The absolute number of SN Ia events simulated and their redshift distribution is calculated using the power-law evolution of the volumetric rates presented by C. Frohmaier et al. (2019) (their Equation 6). The intrinsic distributions of SN stretch and colour are taken from D. Scolnic & R. Kessler (2016).
@@ 3.1.2 Simulation of other types of transients
All simulated transient types are listed in Table LABEL:tab:snClasses. For core-collapse SNe, we use V19 templates and include both hydrogen-rich core-collapse SN events (type II and IIn SNe) and stripped envelope SNe (Type Ib/IIb/Ic/Ic-BL SNe). For volumetric rates of core-collapse events, we normalise at low redshift to the measurement of C. Frohmaier et al. (2021) and extrapolate to higher redshifts assuming the rate follows the star formation history presented in P. Madau & M. Dickinson (2014). To ensure the correct mix of different core collapse sub-types we use the relative rates presented by I. Shivvers et al. (2017). The intrinsic brightness of simulated events are matched to the luminosity functions presented in W. Li et al. (2011) and revised by V19. Recently, updates have been made to models using different rate assumptions for Type IIn SNe (C. L. Ransome & V. A. Villar, 2025) and templates for stripped-envelope (S. Khakpash et al., 2024), however, these updates do not affect the overall picture for TiDES compared to the models we have used.
We include two types of peculiar SNe Ia: SN1991bg-like events (A. V. Filippenko et al., 1992) and SNe Iax (R. J. Foley et al., 2013). SN1991T-like events are modelled within the normal-SN Ia events in Section 3.1.1. The SED templates and volumetric rates used to simulate these classes of peculiar SNe Ia are those presented by R. Kessler et al. (2019a), applying the revisions in M. Vincenzi et al. (2021, hereafter V21).
Tidal Disruption Events (TDEs), Calcium-Strong Transients (CaST) and hydrogen-poor Superluminous SNe (SLSNe-I) are implemented following the same assumptions and modelling used in the PLASTICC challenge (R. Kessler et al., 2019a).
@@ 3.1.3 Simulation of AGN for TiDES-RM
The AGN that we will observe as part of the reverberation mapping survey are already known and are located only in the LSST DDFs. The selection is based on two catalogues, the Million Quasar Catalogue Milliquas (E. W. Flesch, 2023) and the southern photometric quasar catalogue based on Dark Energy Survey DR2 as presented by Q. Yang & Y. Shen (2023). First, both catalogues were cut down to the hexagonal footprints of the 4MOST field-of-view centered on the defined 4MOST deep field positions within the Elais South, COSMOS, ECDF-S, and XMM-LSS deep fields (Table 2). For Milliquas, only sources with a Q or A type flag have been included, which selects against obscured AGN that are unfavourable to reverberation mapping. In a second step, sources from both catalogues were cross-matched. Those present in both catalogues are deemed the “gold sample” while objects only present in Milliquas are part of the “silver sample.” The total number of identified AGN varies with deep field, given how intensely the respective field has been monitored and in which wavebands. Given survey operational constraints, the maximum number of TiDES-RM sources that can be observed per deep field is of the order of 250. To preserve some comparability across the fields, we impose brightness cuts of for the COSMOS and XMM-LSS fields and for the Elais South and ECDF-S fields.
@@
Finally, we want to exclude sources where the survey length will not be able to recover a lag. This requires estimating observed lag length of the cut down catalogues. We use the lag-luminosity relation for the H line presented in M. C. Bentz et al. (2009) and scale to Mg II lags with a factor of 0.9. C IV lags are estimated from the lag-luminosity relation in A. Penton et al. (2022). Approximate luminosities are calculated from band magnitudes. The resulting estimated rest-frame lags have been converted to observed lags by accounting for a cosmological time dilation factor of . We assume that successful lag recovery requires lag lengths to be at least 3 times shorter than the survey length. As such, we cut the target catalogue for AGN where the shortest lag observed in the 4MOST wavelength range for the given redshift is 600 days. With the previously imposed brightness limit, this primarily affects the brightest sources in the redshift range and particularly those at the higher redshift end, where Mg II is the only one of the 3 main target lines present. For all four DDFs, this affects 1% of the pre-cut samples.
| Field | R.A. | Dec. | AGN | H | Mg II | C IV |
| ELAIS-S1 | 9.500 | -43.950 | 191 | 23 | 128 | 80 |
| XMM-LSS | 35.875 | -5.025 | 252 | 47 | 141 | 119 |
| ECDF-S | 53.125 | -28.100 | 230 | 48 | 138 | 92 |
| COSMOS | 150.125 | 2.200 | 256 | 47 | 150 | 94 |
@@
Notes — Redshift ranges covered by individual lines are: H: ; Mg II: ; C IV: . Spectra of some sources will include two broad lines that can be used to cross-calibrate lags: ELAIS-S1: H/Mg II: 10, Mg II/C IV: 30; XMM-LSS: H/Mg II: 17, Mg II/C IV: 38; ECDF-S: H/Mg II: 20, Mg II/C IV: 28; COSMOS: H/Mg II: 14, Mg II/C IV: 21.
@@ 3.2 Simulating transient host galaxies
SNe Ia, peculiar SNe Ia and core-collapse SNe are associated to host galaxies following the approach and assumptions presented in V21. Host galaxies are selected from a ‘HOSTLIB’ library of 380 000 galaxies from the DES Science Verification galaxy catalogues (A. Drlica-Wagner et al., 2018). Each galaxy in the catalogue has optical photometry from DES () and – when available – near infrared photometry from VISTA. To derive properties of the galaxies, template SEDs were generated using the PÉGASE.2 spectral synthesis code and fit to the observed galaxy fluxes through a minimisation; this routine is extensively described in M. Sullivan et al. (2010) and, in the context of DES galaxy catalogues, M. Smith et al. (2020). From the best fitting templates, we adopt the associated parameters such as: redshift, stellar mass, and star formation rates throughout the rest of this work. The catalogue is complete to mag in -band, which is well-matched to the transient hosts that LSST should detect (e.g., P. Wiseman et al., 2020).
Host galaxies are associated to simulated SNe Ia following the observed rates as a function of the host-galaxy properties as measured by M. Sullivan et al. (2006). Within each galaxy, we follow the standard snana prescription of placing transients in the host following a Sérsic profile (J. L. Sérsic, 1963). We also ensure that the well-known correlation between the SN Ia parameter and host galaxy stellar mass is reproduced in the simulations (see section 4.5.1 in V21 for details), but we do not include the more subtle relation between SN Ia Hubble residuals and host galaxy properties, such as stellar mass, colour, and specific star formation rate.
For peculiar SNe Ia, we scale the SN Ia rate and assign host galaxies so that SN1991bg-like events are only associated to old stellar environments (e.g. W. Li et al., 2011; F. H. Panther et al., 2019) and SN Iax events are placed in star forming galaxies (e.g. J. D. Lyman et al., 2013, 2018; T. Takaro et al., 2020). For core-collapse SNe, we use the rates and host galaxy association measurements presented by O. Graur et al. (2017a, b). For the remaining transients their relation to host galaxies is not as extensively studied due to their rarer intrinsic and observational rates; we assign them to hosts randomly and position them within the host following a Sérsic profile. This random host association for the sub-dominant populations has no significant effect on the TiDES host galaxy statistics.
@@ 3.3 Simulating AGN
AGN simulations have been carried out based on properties of the selected sample. Due to the nature of optical selection in the target catalogues, the overwhelming majority of the sources will be broad line AGN, as desired for a reverberation mapping experiment. We identify each selected source with a spectroscopic template of a type 1 AGN as provided by the 4MOST AGN survey (A. Merloni et al., 2019). The associated template spectrum for each object has been redshifted and flux scaled according to the catalogue -band magnitude and fed into the 4MOST operations simulator. Within a 4MOST simulation, the AGN templates are used to determine the required exposure times in the DDFs to meet the spectral success criteria (Section 4.2).
@@ 3.4 Simulating 4MOST
Having generated a simulation of transient light curves as observed by LSST, we next determine which are available to be observed by 4MOST. The 4MOST observing strategy fulfils a complex balance of science requirements from multiple different surveys focused on both galactic and extra-galactic fields. To accommodate this, 4MOST will tile the entire southern-sky over the course of its five-year lifetime (R. S. de Jong et al., 2019), with the density of targets and exposure time requirements ultimately dictating the frequency of any returning visits. The overlap between the LSST wide-fast-deep footprint and the 4MOST union of observing tiles is shown in Fig. 2.
@@
4MOST has experimented with several mock strategies to establish its observing plan. The success of any strategy is evaluated using a figure-of-merit for each 4MOST survey, combined into a global figure of merit. We evaluate the success of TiDES using the total number of objects observed and the quality of their resulting spectra. Our simulations in this paper use a typical 4MOST mock strategy that has been optimised to a solution that tiles the sky with observations (E. Tempel et al., 2020b) using probabilistic fibre assignments (E. Tempel et al., 2020a). From the TiDES perspective, we can use these simulations to track which of our simulated transients were observed and the exposure time on that epoch. For the host galaxies, we are able to assess the total integrated time the target received over multiple epochs.